# Asymptotic behaviors of subcritical branching killed spectrally negative Lévy process

### v1

@@ 3. Conditioned limit theorems for Lévy processes
(3.110) $I_{1}(t)\geq$ $\frac{2R^{*}(x)}{\sqrt{2\pi}\Psi^{\prime\prime}(\lambda_{*})t}\int_{\mathbb{R}_{+}}\left[g(w)1_{\{w\geq\varepsilon\}}-c_{4}\varepsilon h(w)\right]\rho\left(\frac{w}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})t}}\right)\mathrm{d}w$
@@
$J(u):=\int_{0}^{\varepsilon^{1/6}\sqrt{[t]}}\left[\frac{1}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\phi\left(\frac{u-y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\right)\right]\mathbf{P}_{x}^{\lambda_{*}}\left(\xi_{t-m}\in\mathrm{d}y,\tau_{0}^{-}>t-m\right).$
@@
×$\quad\times\left[\mathcal{R}\left(\frac{\varepsilon^{1/6}\sqrt{[t]}}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})(t-m)}}\right)-\mathcal{R}\left(\frac{y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})(t-m)}}\right)\right]\mathrm{d}y.$
Using integration by parts and the fact $F_{u}(0)\geq 0$ (see [7, (3.33)]), we get that there exists a constant $c_{10}$ (independent of $\varepsilon$) such that
$\int_{0}^{\varepsilon^{1/6}\sqrt{[t]}}F_{u}^{\prime}(y)\left[\mathcal{R}\left(\frac{\varepsilon^{1/6}\sqrt{[t]}}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})(t-m)}}\right)-\mathcal{R}\left(\frac{y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})(t-m)}}\right)\right]\mathrm{d}y$
@@
Using [7, Lemma 3.4] with $u=v=\frac{1}{2}\varepsilon^{1/6}\sqrt{[t]}$, since $m=[\varepsilon t]$, we get that there exist positive constants $c_{24}$ and $c_{25}$ both independent of $\varepsilon$ such that
@@
Assume that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1), (H2) and $\mathbf{E}_{0}[\xi_{1}]<0$. Then one can find positive constants $C_{11}$ and $q$ with the property that for any $\varepsilon\in(0,\varepsilon_{0})$ there exist positive constants $T_{4}(\varepsilon)$ and $C_{12}(\varepsilon)$ such that for any $x>0$, $t>T_{4}(\varepsilon)$ and any Borel functions $h,H,g:\mathbb{R}\mapsto\mathbb{R}_{+}$ satisfying $g\leq_{\varepsilon}h\leq_{\varepsilon}H$ and $\int_{\mathbb{R}_{+}}H(z-\varepsilon)(1+z)\mathrm{d}z<\infty$,
@@
Recall that the functions $H_{m}$ and $I_{m}$ are defined in (3.82) and (3.83). Fix $\varepsilon\in(0,\varepsilon_{0})$ and let $h,H,g:\mathbb{R}\mapsto\mathbb{R}_{+}$ be Borel functions satisfying $g\leq_{\varepsilon}h\leq_{\varepsilon}H$ and $\int_{\mathbb{R}_{+}}H(z-\varepsilon)(1+z)\mathrm{d}z<\infty$. For any $y\in\mathbb{R}$, define
@@
Proof of Theorem 3.6: Since $h:\mathbb{R}\mapsto\mathbb{R}_{+}$ is a Borel function and $z\mapsto h(z)(1+|z|)$ is directly Riemann integrable, by [7, Lemma 2.3], there exists $a\in(0,1)$ such that $\int_{\mathbb{R}}\bar{h}_{a,\varepsilon}(1+|z|)\mathrm{d}z<\infty$, for any $\varepsilon\in(0,a)$, where $\bar{h}_{a,\varepsilon}$ is defined in (3.49). Applying Lemma 3.8 to $h$, we have for $t>T_{2}(\varepsilon)$,
@@
Since $h$ is not almost everywhere 00 on $(0,\infty)$, we have
@@
(3.152) $\limsup_{t\to\infty}\frac{t^{3/2}\mathbf{E}_{x}^{\lambda_{*}}\left(h(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right)}{\frac{2R^{*}(x)}{\sqrt{2\pi\Psi^{\prime\prime}(\lambda_{*})^{3}}}\int_{\mathbb{R}_{+}}h(z)\widehat{R}^{*}(z)\mathrm{d}z}$ $\geq\liminf_{t\to\infty}\left[J(\varepsilon,m)-C_{11}\varepsilon I(\varepsilon,m)\right],$
@@
Repeating the argument for $I(y,\varepsilon,m)$ on [7, pp. 40–41], we get
@@
The following result is [9, Lemma 2.1] which is true for any branching killed Lévy process.
@@
(4.4) $u_{h}(x,t):=\mathbb{E}_{x}\left[e^{-\int_{\mathbb{R}_{+}}h(y)X_{t}(\mathrm{d}y)}\right],\quad t>0,~{}x>0,$
@@
(4.5) $u_{h}(x,t)=\mathbf{E}_{x}\left[e^{-h(\xi_{t\land\tau_{0}^{-}})}\right]+\beta\mathbf{E}_{x}\left[\int_{0}^{t}\left(\sum_{k=0}^{\infty}p_{k}u_{h}(\xi_{s\land\tau_{0}^{-}},t-s)^{k}-u_{h}(\xi_{s\land\tau_{0}^{-}},t-s)\right)\mathrm{d}s\right].$
@@
(4.6) $v_{h}(x,t)=\mathbf{E}_{x}\left[1-e^{-h(\xi_{t\land\tau_{0}^{-}})}\right]-\mathbf{E}_{x}\left[\int_{0}^{t}\Phi(v_{h}(\xi_{s\land\tau_{0}^{-}},t-s))\mathrm{d}s\right].$
@@
(4.7) $Q_{y}(x,t)=e^{-\alpha t}\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{0}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right].$
@@
(4.9) $=\lim_{\theta\to\infty}\mathbb{E}_{x}\left(e^{-\int_{\mathbb{R}_{+}}\theta 1_{(y,\infty)}(z)Z^{0}_{t}(\mathrm{d}z)}\right)=:\lim_{\theta\to\infty}u_{\theta 1_{(y,\infty)}}(x,t).$
Now applying Lemma 4.1 we get
$Q_{y}(x,t)$ $=\lim_{\theta\to\infty}\mathbf{E}_{x}\left[1-e^{-\theta 1_{(y,\infty)}(\xi_{t\land\tau_{0}^{-}})}\right]-\mathbf{E}_{x}\left[\int_{0}^{t}\Phi(Q_{y}(\xi_{s\land\tau_{0}^{-}},t-s))\mathrm{d}s\right]$
$=\mathbf{P}_{x}\left(\xi_{t\land\tau_{0}^{-}}>y\right)-\mathbf{E}_{x}\left[\int_{0}^{t}\Phi(Q_{y}(\xi_{s\land\tau_{0}^{-}},t-s))\mathrm{d}s\right].$
This combined with the Feynman-Kac formula gives the desired result. ∎
@@
Assume that (1.7) holds and that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1). Let $x>0$.
@@
    If $\mathbf{E}_{0}\left[\xi_{1}\right]=0$, then for any $y\geq 0$, we have
@@
    If $\mathbf{E}_{0}\left[\xi_{1}\right]>0$, we have
    (4.11) $\liminf_{t\to\infty}e^{\alpha t}u(x,t)\geq C_{sub}\mathbf{E}_{0}\left[\xi_{1}\right]W^{(0)}(x).$
@@
    (4.12) $\liminf_{t\to\infty}e^{\alpha t}Q_{\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t}(x,t)\geq C_{sub}\mathbf{E}_{0}\left[\xi_{1}\right]W^{(0)}(x)\int_{\frac{y}{\sigma}}^{\infty}\phi(z)\mathrm{d}z.$
@@
    Assume further that (H2) holds. If $\mathbf{E}_{0}\left[\xi_{1}\right]<0$, then for any $y\geq 0$, we have
@@
(4.15) $Q_{y}(x,t)=e^{-\alpha t}\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{0}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right]$
(4.16) $\geq e^{-\alpha t}e^{-\int_{0}^{t}\varphi(g(t-s))\mathrm{d}s}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}>y\right)\geq C_{sub}e^{-\alpha t}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}>y\right).$
Applying Lemma 3.1, we immediately get the assertion of (1). Using the fact that $u(x,t)=Q_{0}(x,t)$ and applying Theorem 2.1, we get the first result of (2). Applying Lemma 2.3 with $y$ replaced by $\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t$, we get the second result of (2). Applying Theorem 3.5 with $f(x)=1_{(y,\infty)}(x)$, we get the assertion of (3). ∎
@@
Lemma 4.4.
Assume that (1.7) holds and that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1). If $\mathbf{E}_{0}\left[\xi_{1}\right]=0$, then for any $x>0$ and $y\geq 0$, we have
@@
Fix an $N>0$. For $t>N$, by Lemma 4.2, we have
(4.18) $Q_{\sqrt{t}y}(x,t)$ $\leq e^{-\alpha t}\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y\}}e^{-\int_{t-N}^{t}\varphi(Q_{\sqrt{t}y}(\xi_{s},t-s))\mathrm{d}s}\right]$
(4.19) $\leq e^{-\alpha t}\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y\}}e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y}(\inf_{r\in[t-N,t]}\xi_{r},s))\mathrm{d}s}\right].$
@@
(4.20) $J_{1}(t):=\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y,\inf_{r\in[t-N,t]}\xi_{r}\geq\sqrt{t}y+t^{\gamma}\}}e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y}(\inf_{r\in[t-N,t]}\xi_{r},s))\mathrm{d}s}\right],$
(4.21) $J_{2}(t):=\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+t^{\gamma}\}}e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y}(\inf_{r\in[t-N,t]}\xi_{r},s))\mathrm{d}s}\right].$
@@
(4.24) $e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y}(\sqrt{t}y+t^{\gamma},s))\mathrm{d}s}\leq\exp\Big{\{}-\int_{0}^{N}\varphi\left(g(s)\mathbf{P}_{\sqrt{t}y+t^{\gamma}}\left(\tau_{0}^{-}>s,\xi_{s}>\sqrt{t}y\right)\right)\Big{\}}.$
@@
[…]
@@
where $W$ is the Brownian motion in Lemma 3.2. Then by the Markov property of $\xi$, for $k<t-N$,
@@
(4.33) $H_{1}(t)$ $\leq\frac{CN_{*}(\kappa)}{(t-k)^{(\frac{1}{2}-\kappa)(\delta+2)-1}}\frac{1+x}{\sqrt{k}},$
@@
Assume that (1.7) holds and that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1). If $\mathbf{E}_{0}\left[\xi_{1}\right]>0$, then for any $x>0$, we have
(4.39) $\limsup_{t\to\infty}e^{\alpha t}u(x,t)\leq C_{sub}\mathbf{E}_{0}\left[\xi_{1}\right]W^{(0)}(x).$
@@
(4.40) $\limsup_{t\to\infty}e^{\alpha t}Q_{\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t}(x,t)\leq C_{sub}\mathbf{E}_{0}\left[\xi_{1}\right]W^{(0)}(x)\int_{\frac{y}{\sigma}}^{\infty}\phi(z)\mathrm{d}z.$

### v2

@@ 3. Conditioned limit theorems for Lévy processes
(3.110) $I_{1}(t)\geq$ $\frac{2R^{*}(x)}{\sqrt{2\pi}\Psi^{\prime\prime}(\lambda_{*})t}\int_{\mathbb{R}_{+}}\left(g(w)1_{\{w\geq\varepsilon\}}-c_{4}\varepsilon h(w)\right)\rho\left(\frac{w}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})t}}\right)\mathrm{d}w$
@@
$J(u):=\int_{0}^{\varepsilon^{1/6}\sqrt{[t]}}\left(\frac{1}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\phi\left(\frac{u-y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\right)\right)\mathbf{P}_{x}^{\lambda_{*}}\left(\xi_{t-m}\in\mathrm{d}y,\tau_{0}^{-}>t-m\right).$
@@
$\quad\times\left(\mathcal{R}\left(\frac{\varepsilon^{1/6}\sqrt{[t]}}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})(t-m)}}\right)-\mathcal{R}\left(\frac{y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})(t-m)}}\right)\right)\mathrm{d}y.$
Using integration by parts and the fact $F_{u}(0)\geq 0$ (see [8, (3.33)]), we get that there exists a constant $c_{10}$ (independent of $\varepsilon$) such that
$\int_{0}^{\varepsilon^{1/6}\sqrt{[t]}}F_{u}^{\prime}(y)\left(\mathcal{R}\left(\frac{\varepsilon^{1/6}\sqrt{[t]}}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})(t-m)}}\right)-\mathcal{R}\left(\frac{y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})(t-m)}}\right)\right)\mathrm{d}y$
@@
Using [8, Lemma 3.4] with $u=v=\frac{1}{2}\varepsilon^{1/6}\sqrt{[t]}$, since $m=[\varepsilon t]$, we get that there exist positive constants $c_{24}$ and $c_{25}$ both independent of $\varepsilon$ such that
@@
Assume that $\xi$ is a Lévy process satisfying (H1), (H2) (H3) and $\mathbf{E}_{0}[\xi_{1}]<0$. Then one can find positive constants $C_{11}$ and $q$ with the property that for any $\varepsilon\in(0,\varepsilon_{0})$ there exist positive constants $T_{4}(\varepsilon)$ and $C_{12}(\varepsilon)$ such that for any $x>0$, $t>T_{4}(\varepsilon)$ and any Borel functions $h,H,g:\mathbb{R}\to\mathbb{R}_{+}$ satisfying $g\leq_{\varepsilon}h\leq_{\varepsilon}H$ and $\int_{\mathbb{R}_{+}}H(z-\varepsilon)(1+z)\mathrm{d}z<\infty$,
@@
Recall that the functions $H_{m}$ and $I_{m}$ are defined in (3.82) and (3.83). Fix $\varepsilon\in(0,\varepsilon_{0})$ and let $h,H,g:\mathbb{R}\to\mathbb{R}_{+}$ be Borel functions satisfying $g\leq_{\varepsilon}h\leq_{\varepsilon}H$ and $\int_{\mathbb{R}_{+}}H(z-\varepsilon)(1+z)\mathrm{d}z<\infty$. For any $y\in\mathbb{R}$, define
@@
Proof of Theorem 3.6: Since $h:\mathbb{R}\to\mathbb{R}_{+}$ is a Borel function and $z\mapsto h(z)(1+|z|)$ is directly Riemann integrable, by [8, Lemma 2.3], there exists $a\in(0,1)$ such that $\int_{\mathbb{R}}\bar{h}_{a,\varepsilon}(1+|z|)\mathrm{d}z<\infty$, for any $\varepsilon\in(0,a)$, where $\bar{h}_{a,\varepsilon}$ is defined in (3.49). Applying Lemma 3.8 to $h$, we have for $t>T_{2}(\varepsilon)$,
@@
Since $h$ is not almost everywhere 0 on $(0,\infty)$, we have
@@
(3.152) $\limsup_{t\to\infty}\frac{t^{3/2}\mathbf{E}_{x}^{\lambda_{*}}\left(h(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right)}{\frac{2R^{*}(x)}{\sqrt{2\pi\Psi^{\prime\prime}(\lambda_{*})^{3}}}\int_{\mathbb{R}_{+}}h(z)\widehat{R}^{*}(z)\mathrm{d}z}$ $\geq\limsup_{t\to\infty}\left(J(\varepsilon,m)-C_{11}\varepsilon I(\varepsilon,m)\right),$
@@
Repeating the argument for $I(y,\varepsilon,m)$ on [8, pp. 40–41], we get
@@
Let $B_{b}^{+}(\mathbb{R}_{+})$ be the space of non-negative bounded Borel functions on $\mathbb{R}_{+}$. The following result is [11, Lemma 2.1] which is true for any branching killed Lévy process.
@@
(4.4) $u_{h}(x,t):=\mathbb{E}_{x}\left(e^{-\int_{\mathbb{R}_{+}}h(y)Z_{t}^{0}(\mathrm{d}y)}\right),\quad t>0,~x>0,$
@@
(4.5) $u_{h}(x,t)=\mathbf{E}_{x}\left(e^{-h(\xi_{t\land\tau_{0}^{-}})}\right)+\beta\mathbf{E}_{x}\left(\int_{0}^{t}\left(\sum_{k=0}^{\infty}p_{k}u_{h}(\xi_{s\land\tau_{0}^{-}},t-s)^{k}-u_{h}(\xi_{s\land\tau_{0}^{-}},t-s)\right)\mathrm{d}s\right).$
@@
(4.6) $v_{h}(x,t)=\mathbf{E}_{x}\left(1-e^{-h(\xi_{t\land\tau_{0}^{-}})}\right)-\mathbf{E}_{x}\left(\int_{0}^{t}\Phi(v_{h}(\xi_{s\land\tau_{0}^{-}},t-s))\mathrm{d}s\right).$
@@
(4.7) $Q_{y}(x,t)=e^{-\alpha t}\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{0}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right).$
@@
(4.9) $=\lim_{\theta\to\infty}\mathbb{E}_{x}\left(e^{-\int_{\mathbb{R}_{+}}\theta 1_{(y,\infty)}(z)Z^{0}_{t}(\mathrm{d}z)}\right).$
Now applying Lemma 4.1 with $h(z)=1_{(y,\infty)}(z)$, we get
(4.10) $Q_{y}(x,t)$ $=\lim_{\theta\to\infty}\mathbf{E}_{x}\left(1-e^{-\theta 1_{(y,\infty)}(\xi_{t\land\tau_{0}^{-}})}\right)-\mathbf{E}_{x}\left(\int_{0}^{t}\Phi(Q_{y}(\xi_{s\land\tau_{0}^{-}},t-s))\mathrm{d}s\right)$
(4.11) $=\mathbf{P}_{x}\left(\xi_{t\land\tau_{0}^{-}}>y\right)-\mathbf{E}_{x}\left(\int_{0}^{t}\Phi(Q_{y}(\xi_{s\land\tau_{0}^{-}},t-s))\mathrm{d}s\right).$
Thus $Q_{y}(x,t)$ is a bounded solution of the following equation
(4.12) $u(x,t)=\mathbf{P}_{x}\left(\xi_{t\land\tau_{0}^{-}}>y\right)-\mathbf{E}_{x}\left(\int_{0}^{t}\Phi(u(\xi_{s\land\tau_{0}^{-}},t-s))\mathrm{d}s\right).$
It follows from [15, (4.8), p.102] that there is a unique positive locally bounded solution to (4.12). Thus we only need to prove that the right side of (4.7) is also a solution (4.12). For $s\in[0,t]$, define
(4.13) $A_{s,t}=-\int_{s}^{t}\frac{\Phi(Q_{y}(\xi_{r\land\tau_{0}^{-}},t-r))}{Q_{y}(\xi_{r\land\tau_{0}^{-}},t-r)}\mathrm{d}r.$
Note that $\frac{\Phi(u)}{u}=\varphi(u)+\alpha$ for $u\in(0,1]$. The right side of (4.7) can be written as $\mathbf{E}_{x}\left(e^{A_{0,t}}1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}\right)$. It is elementary to check that
(4.14) $e^{A_{0,t}}=1-\int_{0}^{t}e^{A_{s,t}}\frac{\Phi(Q_{y}(\xi_{s\land\tau_{0}^{-}},t-s))}{Q_{y}(\xi_{s\land\tau_{0}^{-}},t-s)}\mathrm{d}s.$
Hence we have
(4.15) $\mathbf{E}_{x}\left(e^{A_{0,t}}1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}\right)=$ $\mathbf{P}_{x}\left(\xi_{t\land\tau_{0}^{-}}>y\right)$
(4.16) - $\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}\int_{0}^{t}e^{A_{s,t}}\frac{\Phi(Q_{y}(\xi_{s},t-s))}{Q_{y}(\xi_{s},t-s)}\mathrm{d}s\right).$
Now using the Markov property and the fact that
(4.17) $A_{s,t}=\int_{0}^{t-s}\frac{\Phi(Q_{y}(\xi_{(r+s)\land\tau_{0}^{-}},t-r-s))}{Q_{y}(\xi_{(r+s)\land\tau_{0}^{-}},t-r-s)}\mathrm{d}r,$
we see that (4.15) implies that $\mathbf{E}_{x}\left(e^{A_{0,t}}1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}\right)$ solves (4.12). Thus, we have
(4.18) $Q_{y}(x,t)$ $=\mathbf{E}_{x}\left(e^{A_{0,t}}1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}\right)$
(4.19) $=e^{-\alpha t}\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{0}^{t}\varphi(Q_{y}(\xi_{r},t-s))\mathrm{d}s}\right).$
This gives the desired result. ∎
@@
Assume that (1.7) holds and $\xi$ is a Lévy process satisfying (H1). Let $x>0$.
@@
    If $\mathbf{E}_{0}\left(\xi_{1}\right)=0$, then for any $y\geq 0$, we have
@@
    If $\mathbf{E}_{0}\left(\xi_{1}\right)>0$, we have
    (4.21) $\liminf_{t\to\infty}e^{\alpha t}u(x,t)\geq q_{x}C_{sub}.$
@@
    (4.22) $\liminf_{t\to\infty}e^{\alpha t}Q_{\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t}(x,t)\geq q_{x}C_{sub}\int_{\frac{y}{\sigma}}^{\infty}\phi(z)\mathrm{d}z.$
@@
    If $\mathbf{E}_{0}\left(\xi_{1}\right)<0$ and (H2) and (H3) hold, then for any $y\geq 0$, we have
@@
(4.25) $Q_{y}(x,t)=e^{-\alpha t}\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>y\}}e^{-\int_{0}^{t}\varphi(Q_{y}(\xi_{s},t-s))\mathrm{d}s}\right)$
(4.26) $\geq e^{-\alpha t}e^{-\int_{0}^{t}\varphi(g(t-s))\mathrm{d}s}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}>y\right)\geq C_{sub}e^{-\alpha t}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}>y\right),$
where the last inequality follows from (1.13). Applying Lemma 3.1, we immediately get the assertion of (1). Using the fact that $u(x,t)=Q_{0}(x,t)$, $\Psi^{\prime}(0+)=\mathbb{E}_{0}(\xi_{1})$ and applying (2.10) and (4.25), we get
(4.27) $\liminf_{t\to\infty}e^{\alpha t}u(x,t)\geq\liminf_{t\to\infty}C_{sub}\mathbf{P}_{x}\left(\tau_{0}^{-}>t\right)=q_{x}C_{sub}.$
This gives the first result of (2). Applying Lemma 2.4 with $y$ replaced by $\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t$, we get the second result of (2). Applying Theorem 3.5 with $f(x)=1_{(y,\infty)}(x)$, we get the assertion of (3). ∎
@@
Lemma $4.$
Assume that (1.7) holds and $\xi$ is a Lévy process satisfying (H1). If $\mathbf{E}_{0}\left(\xi_{1}\right)=0$, then for any $x>0$ and $y\geq 0$, we have
@@
Recall that $\varphi$ and $Q_{y}(\cdot,t)$ are increasing functions. Fix an $N>0$. For $t>N$, by Lemma 4.2, we have
(4.29) $Q_{\sqrt{t}y}(x,t)$ $\leq e^{-\alpha t}\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y\}}e^{-\int_{t-N}^{t}\varphi(Q_{\sqrt{t}y}(\xi_{s},t-s))\mathrm{d}s}\right)$
(4.30) $\leq e^{-\alpha t}\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y\}}e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y}(\inf_{r\in[t-N,t]}\xi_{r},s))\mathrm{d}s}\right).$
@@
(4.31) $J_{1}(t):=\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y,\inf_{r\in[t-N,t]}\xi_{r}\geq\sqrt{t}y+t^{\gamma}\}}e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y}(\inf_{r\in[t-N,t]}\xi_{r},s))\mathrm{d}s}\right),$
(4.32) $J_{2}(t):=\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y,\inf_{r\in[t-N,t]}\xi_{r}<\sqrt{t}y+t^{\gamma}\}}e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y}(\inf_{r\in[t-N,t]}\xi_{r},s))\mathrm{d}s}\right).$
@@
(4.35) $e^{-\int_{0}^{N}\varphi(Q_{\sqrt{t}y}(\sqrt{t}y+t^{\gamma},s))\mathrm{d}s}\leq\exp\Big\{-\int_{0}^{N}\varphi\left(g(s)\mathbf{P}_{\sqrt{t}y+t^{\gamma}}\left(\tau_{0}^{-}>s,\xi_{s}>\sqrt{t}y\right)\right)\mathrm{d}s\Big\}.$
@@
(4.36) $\limsup_{N\to\infty}\limsup_{t\to\infty}\frac{J_{1}(t)}{\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}>\sqrt{t}y\right)}\leq\limsup_{N\to\infty}e^{-\int_{0}^{N}\varphi(g(s))\mathrm{d}s}=C_{sub}.$
@@
where $W$ is the Brownian motion in Lemma 3.2. Then by the Markov property of $\xi$, for k<tNk<t-N,
@@
(4.44) $H_{1}(t)$ $\leq\frac{CC_{3}(\kappa)}{(t-k)^{(\frac{1}{2}-\kappa)(\delta+2)-1}}\frac{1+x}{\sqrt{k}},$
@@
Assume that (1.7) holds and that $\xi$ is a Lévy process satisfying (H1). If $\mathbf{E}_{0}\left(\xi_{1}\right)>0$, then for any $x>0$, we have
(4.50) $\limsup_{t\to\infty}e^{\alpha t}u(x,t)\leq q_{x}C_{sub}.$
@@
(4.51) $\limsup_{t\to\infty}e^{\alpha t}Q_{\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t}(x,t)\leq q_{x}C_{sub}\int_{\frac{y}{\sigma}}^{\infty}\phi(z)\mathrm{d}z.$